# Fine-Tuning Large Language Models to Translate: Will a Touch of Noisy Data in Misaligned Languages Suffice?

### v1

@@ Abstract
Traditionally, success in multilingual machine translation can be attributed to three key factors in training data: large volume, diverse translation directions, and high quality. In the current practice of fine-tuning large language models (LLMs) for translation, we revisit the importance of all these factors. We find that LLMs display strong translation capability after being fine-tuned on as few as 32 training instances, and that fine-tuning on a single translation direction effectively enables LLMs to translate in multiple directions. However, the choice of direction is critical: fine-tuning LLMs with English on the target side can lead to task misinterpretation, which hinders translations into non-English languages. A similar problem arises when noise is introduced into the target side of parallel data, especially when the target language is well-represented in the LLM’s pre-training. In contrast, noise in an under-represented language has a less pronounced effect. Our findings suggest that attaining successful alignment hinges on teaching the model to maintain a “superficial” focus, thereby avoiding the learning of erroneous biases beyond translation.
@@
Dawei Zhu1 Pinzhen Chen2 Miaoran Zhang1 Barry Haddow2 Xiaoyu Shen3 Dietrich Klakow1 1Saarland University, Saarland Informatics Campus 2University of Edinburgh 3Eastern Institute of Technology {dzhu, mzhang}@lsv.uni-saarland.de pinzhen.chen@ed.ac.uk
@@ 1 Introduction
Large language models (LLMs) have reached heights in various NLP tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023). Supervised fine-tuning (SFT (Ouyang et al., 2022), also called instruction tuning or simply fine-tuning in some literature) further prepares these models for better generalization and reliability in downstream tasks by training on input-output data combined with task instructions in natural languages (Sanh et al., 2022; Wei et al., 2022; Mishra et al., 2022). In this research direction, various works have studied the “scaling up” of data size, number of languages, etc, and demonstrated strong results (Chung et al., 2022; Muennighoff et al., 2023). On the other hand, more recent papers also embraced the philosophy of “less is more” by curating a small set of high-quality training instances, claiming a “superficial alignment hypothesis” (Zhou et al., 2023) with similar findings by others (e.g. Chen et al., 2024a).
This work further investigates the role of SFT data in preparing LLMs for machine translation (MT), a cross-lingual generation task with high demands in practical domains. Prior research has found fine-tuning to improve translation performance (Zhang et al., 2023c) and more recent works also integrated continued pre-training with more data to provide further improvement (Xu et al., 2024a; Alves et al., 2024). Nonetheless, the feasibility of “less is more” in translation is rather under-explored. In translation prompting, researchers have suggested that a model’s translation capability can be attributed to the bilingual signals exposed during pre-training (Briakou et al., 2023) and task recognition in LLM layers (Sia et al., 2024), hinting that the translation task has been picked up during pre-training. A natural question follows: can we put reduced effort into data?
From a practical perspective, we squeeze the translation SFT data to a mere size of 32 or the translation direction to 1 for multilingual translation, for which we believe LLMs already possess a strong pre-trained foundation in multilingual understanding and generation. Beyond quantity and language diversity, we perform SFT on data with inferior quality which could be common for lower-resourced languages. To summarize, our analysis is grounded in the task of MT, with “scaling down” in mind. In multiple dimensions—data size (§3.2), translation direction (§3.3, §3.4), and data quality (§3.5)—our findings verify, complement, and refine the existing superficial alignment hypothesis:
@@
32 data instances successfully enable an LLM to translate in 11 directions. More data still helps but the returns diminish.
@@
Data in a single translation direction can effectively align an LLM to translate to and from multiple directions. Yet, it is crucial to pick the right direction—we recommend not to place English on the target side.
@@
Injecting noise into the SFT data results in different patterns concerning language exposure. An LLM can easily overfit to the noise patterns in high-resource languages while it is more robust to data noise in low-resource languages.
@@ 2.2 Superficial alignment hypothesis
Zhou et al. (2023) claim that a model’s knowledge and capabilities are acquired almost entirely during pre-training, and the effect of alignment tuning might be “superficial”, in that it teaches the model the format for interacting with users. This idea is further supported by recent works (Lin et al., 2024; Ghosh et al., 2024). However, to what extent this applies to multilingual translation remains underexplored. To bridge this gap, We conduct a series of controlled experiments to fine-tune LLMs for translation, complementing previous research across three dimensions. First, we study the parallel data efficiency in the era of LLMs, aiming to determine the minimum data needed for effective model alignment. Second, we explore the scope of alignment by probing whether aligning one translation direction influences other directions. Finally, we investigate how varying data quality affects the LLMs’ behavior in aligning translation tasks.
@@ 3.1 Experimental setup
By default, we take the test sets from WMT17 to WMT20 as our parallel training data (Bojar et al., 2017, 2018; Barrault et al., 2019, 2020); we also use the development sets in WMT21 (Akhbardeh et al., 2021) for training if a language pair of interest is not available in earlier years. The specific training data configurations will be detailed in the individual subsequent sections. The test sets from WMT21 are used for validation. Detailed data statistics can be found in Section F.1. The LLM we use for SFT is the base version of Llama-2 7B (Touvron et al., 2023). When performing SFT, we use a learning rate of 5e-6, an effective batch size of 64, and linear learning rate scheduling with a warmup ratio of 0.1. We select the model checkpoint based on COMET scores on the validation sets.111We find that the validation loss has a relatively weak correlation with COMET scores measured on the validation set, similar to findings from (Ouyang et al., 2022). To form the model input for SFT, we feed the source sentence into the Alpaca prompt template (Taori et al., 2023), supplementing it with a translation instruction that is randomly selected from a pool of 31 instructions. Refer to Table 4 in the appendix for the complete list of instructions.
@@
We primarily evaluate the models on the WMT22 test sets (Kocmi et al., 2022) covering 11 translation directions: encs, ende, enjp, enru, enzh, and enhr.222Language codes: cs=Czech, de=German, hr=Croatian, jp=Japanese, ru=Russian, zh=Chinese. “” means that both translation directions are covered. Note that only enhr is available in WMT22 but not the inverse direction. Languages in these 11 directions are explicitly included in Llama-2’s pre-training corpus. In Section 3.4, we extend our evaluation to translation directions involving middle and low resource languages: Icelandic and Hausa (i.e., enis, enha), and this data comes from WMT21’s test set. At inference time, a fixed translation instruction is applied to all test examples. We set the maximum generation length to 256 tokens and use beam search with a beam size of 4 for decoding. We use reference-based COMET-22 (Rei et al., 2020)333Specifically, COMET is reported on a scale of 0 to 100 as opposed to its raw 0 to 1 range. and BLEU (Papineni et al., 2002) as the evaluation metrics. Refer to detailed software versions in Section F.3.
3.2 How much training data is required to enable LLMs for translation?
@@
We fine-tune Llama-2 7B using different numbers of training samples and evaluate the multilingual translation performance of the resulting models. We collect training data covering 10 translation directions: en{cs, de, jp, ru, zh}. The training data sourced from WMT17-20 contains a total of 74,623 parallel examples. Note that the training samples across translation directions are not evenly distributed. To create training sets of varying sizes, we subsample the original data into subsets that are powers of 2, starting from 32 () and ending with 4096 (); larger subsets always contain smaller subsets. To ensure balanced language representation in our subsets, we distribute samples as evenly as possible among the language pairs.444For example, the data size distribution for our 32-example training set is . We refer to the fine-tuned model as SFT-MT. Considering LLMs are capable of performing translation through prompting too, we compare SFT-MT with 1- and 3-shot in-context learning (ICL), denoted as ICL-MT. For ICL, we randomly select demonstrations from the training set in the test direction for each test sentence. We do not consider Llama-2’s zero-shot performance because, although it sometimes produces acceptable translations at the beginning, it often does not stop generating, which makes it difficult to accurately evaluate its performance. Lastly, as demonstrated in (Zhu et al., 2024), established LLM-based chatbots trained on diverse NLP tasks also serve as strong translation systems. Therefore, we compare our models with general-purpose, instruction-tuned LLMs (IT-LLM), including Vicuna-v1.5-7b (Chiang et al., 2023), Mistral-7b (Jiang et al., 2023), and Llama-2-7b-chat (Touvron et al., 2023).555lmsys/vicuna-7b-v1.5, mistralai/Mistral-7B-v0.1, and meta-llama/Llama-2-7b-chat-hf.
@@ Results.
Figure 1 illustrates the effect of varying training sample sizes on translation performance. In both 1- and 3-shot cases, ICL-MT underperforms IT-LLM baselines like Llama-2-7b-chat despite sharing the same foundation model, indicating that a few in-context demonstrations may not effectively align Llama-2 for translation. However, performance significantly improves as soon as Llama-2 is fine-tuned with just 32 samples, outperforming all three IT-LLM baselines in both COMET and BLEU metrics. This suggests that a handful of high-quality parallel data can effectively specialize the model into a performant translation system. Increasing parallel data further boosts performance, though with diminishing returns: the COMET score rises by an average of 2 points when expanding from 32 to 1024 samples, but only by 0.5 points when increasing further from 1024 to 75K samples (the full training set). Given that it is unlikely that these 32 training samples “teach” Llama-2 new translation skills, this shows strong evidence that superficial alignment applies to MT. In general, we find that around 1K samples are a sweet spot for data efficiency. In summary, effective translation alignment begins with minimal training data, revealing less is good alignment and more is better with diminishing gains.
@@ 3.3 Do we need to include all directions?
In the preceding section, we follow the traditional practice in multilingual MT by including multiple translation directions during training. However, the observation that Llama-2 may efficiently adapt to translation tasks with only a few dozen samples leads us to reconsider the necessity of including samples from all these directions. Specifically, could training with examples from just a single translation direction be sufficient to align LLMs for multilingual translation?
@@
Figure 2 demonstrates the normalized performance of Llama-2 when fine-tuned in various single directions. Remarkably, training with just one direction enables Llama-2 to translate between multiple languages. For instance, after fine-tuning on deen or zhen, the model can translate from all considered languages to English with at least efficacy relative to training on all directions. Similarly, the model fine-tuned on ende, enzh, frde or defr also demonstrate only a slight performance decline when translating from English. Notable declines are observed in two scenarios: (1) trained to translate to English and evaluated on translating to non-English; and (2) trained to translate to non-English and evaluated on translating to English.666Analysis of model outputs reveals that they often merely echo the source sentence, ignoring the translation instruction. Of these two scenarios, scenario 1 exhibits a much larger performance drop. The fact that both scenarios involve a mismatch between using English and non-English suggests that Llama-2, as an English-centric LLM, may process English differently compared to other languages. When fine-tuned for English generation, the model may misinterpret the task as only generating in English. Generalization among non-English languages is much easier than generalization between English and non-English languages, as evidenced by the negligible performance drop when fine-tuning and testing on two vastly different language pairs such as defr and enzh. Overall, the findings suggest that SFT in one translation direction effectively enables the many directions, though avoiding misinterpretation is crucial.
@@
The results of performing ICL with misaligned directions are presented in Table 1. It can be seen that misaligned demonstrations significantly degrade translation performance, with 3-shot be often worse than 1-shot. We observe that the model may output Chinese characters, emojis, time and date etc., but no clear error patterns are observed. This contrasts sharply with findings from SFT: while SFT can recognize the format of translation, ICL requires language-aligned demonstrations.
@@
Figure 3 presents a joint evaluation on sizes and directions. For small training sizes, covering diverse translation directions in training proves to be beneficial. However, the benefits of such diversity level off as the training size increases. With a training size of 1024, models trained exclusively on two directions, ende, perform on par with those trained on all directions.
3.4 Can alignment be achieved with unseen languages?
Previous sections focus on translation directions involving languages explicitly included in Llama-2’s pre-training corpus. We now extend our investigation to languages that do not have an identified presence of over 0.005% in the pre-training data (c.f. Touvron et al., 2023, p22), referred to as unseen languages. Here we seek answers to two questions: (1) Can we effectively align Llama-2 to translate both from and to unseen languages by fine-tuning it with a small dataset? (2) How well can this fine-tuned model translate from and to seen languages?
@@ Results.
The results are presented in Figure 4. It can be seen that fine-tuning on Icelandic and Hausa enhances a model’s translation quality on these languages compared to the control setup, yet the gains are modest. We observe that Llama-2 manages to produce tokens in these languages, however, the translations often largely deviate from the original meanings. This suggests that it is difficult to teach models new translation directions via SFT with limited data. Interestingly, we find fine-tuning on Icelandic or Hausa does not hinder Llama-2’s ability to translate from and to all seen languages, maintaining performance levels comparable to the control scenario with ende. Based on these results, we propose a complement to the superficial alignment hypothesis in MT: LLMs may learn the essence of the task without requiring a deep understanding of the input-output mapping of the training samples.
3.5 Can we use noisy data?
So far we have observed that LLMs quickly recognize the translation task with minimal high-quality, manually curated data. But what if the data contains noise? Can LLMs still recognize the essence of translation—accurately conveying the meaning of the source sentence in another language—–or will they instead overfit to the noise, leading to degraded translation performance?
Setup.
To create noisy training sets, we inject two types of noise: sentence-level and word-level, into the clean training data. Both types of noise are introduced via back-translation (Sennrich et al., 2016). Specifically, we use the OPUS-MT suite (Tiedemann and Thottingal, 2020) to translate from English to a target non-English language.777For instance, when training on noised deen, OPUS-MT is used to translate from English data to German, introducing noise on the source side. Conversely, for noised ende training, noise is added to the target side. For word-level noising, we translate individual source words by feeding a single space-delimited word at a time. The OPUS-MT checkpoints can be found under Helsinki-NLP/opus-mt-$en-${trg} on Hugging Face. We investigate the impact of noise in four translation directions: ende′, deen, enha′, and haen, where the prime (′) notation indicates the side receiving noise. We consider two sizes of the training sets: 32 and 1024.
Results.
The results are presented in Figure 5. It can be seen that both types of noise generally cause a decline in performance. However, The degree of degradation significantly varies depending on whether the noise is injected into the source or target side of the translation, as well as the language affected. Specifically, when noise is introduced to the target side, models fine-tuned on ende′ and enha′ translations exhibit a clear drop in performance. The impact of word noise is more severe than that of sentence noise. In the case of ende′, word-level noise makes the model degenerate to producing literal translations across all test cases. An example of this behavior is presented in Section 3.5. In contrast, the performance drop caused by word noise is less pronounced with enha′, particularly when evaluated on enX.
Summarizing the observations, Llama-2 is much more robust against the noise added to Hausa, likely because it has limited familiarity with the language, making it more difficult to imitate the noise patterns. As a result, Llama-2 tends to just recognize the task instead of fitting the noise. In contrast, for German, Llama-2’s understanding leads to a misinterpretation of the training objectives, such as fitting the word-level noise with a directive for literal translations. Overall, LLMs may quickly fit translation noise and imitate the noise patterns, especially for seen languages; the resulting performance drop may be observable with just 32 training samples.
| Source | Ref./Training config. | Model output |
@@
| sehr ärgerlich. | Word-to-word ref. | The find I honest said very annoying. |
@@
| Word-to-word ref. | in order to avoid again happen such thing. | |

### v2

@@ Abstract
Traditionally, success in multilingual machine translation can be attributed to three key factors in training data: large volume, diverse translation directions, and high quality. In the current practice of fine-tuning large language models (LLMs) for translation, we revisit the importance of these factors. We find that LLMs display strong translation capability after being fine-tuned on as few as 32 parallel sentences and that fine-tuning on a single translation direction enables translation in multiple directions. However, the choice of direction is critical: fine-tuning LLMs with only English on the target side can lead to task misinterpretation, which hinders translation into non-English languages. Problems also arise when noisy synthetic data is placed on the target side, especially when the target language is well-represented in LLM pre-training. Yet interestingly, synthesized data in an under-represented language has a less pronounced effect. Our findings suggest that when adapting LLMs to translation, the requirement on data quantity can be eased but careful considerations are still crucial to prevent an LLM from exploiting unintended data biases.111Code available at: github.com/uds-lsv/mt-sft.
@@
Dawei Zhu1 Pinzhen Chen2 Miaoran Zhang1 Barry Haddow2 Xiaoyu Shen3††thanks: Corresponding author (xyshen@eitech.edu.cn) Dietrich Klakow1 1Saarland University, Saarland Informatics Campus 2University of Edinburgh 3Digital Twin Institute, Eastern Institute of Technology, Ningbo {dzhu,mzhang}@lsv.uni-saarland.de pinzhen.chen@ed.ac.uk
@@ 1 Introduction
Large language models (LLMs) have reached new heights in various NLP tasks (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Jiang et al., 2023). Supervised fine-tuning (SFT, Ouyang et al., 2022, alternatively, instruction tuning or simply fine-tuning in some literature) further prepares these models for better generalization and reliability in downstream tasks by training on task input-output data combined with instructions in natural languages (Sanh et al., 2022; Wei et al., 2022; Mishra et al., 2022). In this research direction, various works have studied the “scaling up” of SFT data size, number of languages, etc (Chung et al., 2024; Muennighoff et al., 2023). On the other hand, recent papers also embraced the philosophy of “less is more” by achieving strong results with a small set of high-quality training instances, claiming a “superficial alignment hypothesis” (Zhou et al., 2023) with similar findings by others.
This work investigates the role of SFT data in aligning LLMs to machine translation (MT), a cross-lingual generation task with high demands in practical domains. Prior research has found fine-tuning to improve translation performance (Zhang et al., 2023c) and more recent works also integrated continued pre-training with more data to provide further improvement (Xu et al., 2024a; Alves et al., 2024). For encoder-decoder models, Wu et al. (2024a) used little data to enable an English-centric model to translate between any two languages. Nonetheless, the feasibility of “less is more” in LLM translation fine-tuning is rather under-explored. In translation prompting, researchers have suggested that a model’s translation capability can be attributed to the bilingual signals exposed during pre-training (Briakou et al., 2023) and task recognition in LLM layers (Sia et al., 2024), hinting that the translation capability has been picked up during pre-training. A natural question follows: Can we put reduced effort into data?
From a data efficiency perspective, we squeeze the translation SFT data to a mere size of 32 or the translation direction to 1 for multilingual translation, for which we believe LLMs already possess a strong pre-trained foundation in multilingual understanding and generation. Beyond quantity and language diversity, we perform SFT on synthesized data via machine translation, which is a common data augmentation practice for under-served languages. To summarize, our analysis is grounded in the task of MT, with “scaling down” in mind. In multiple dimensions—data size (§3.2), translation direction (§3.3 and §3.4), and data synthesis (§3.5)—our findings verify, complement, and refine the existing superficial alignment hypothesis for fine-tuning LLMs for translation tasks:
@@
32 data instances successfully enable an LLM to translate in 11 directions. More data still helps but the return diminishes.
@@
Data in a single translation direction can effectively align an LLM to translate to and from multiple directions. Yet, it is crucial to pick the right direction—we recommend not placing English on the target side.
@@
When fine-tuning on lower-quality synthetic data, LLMs are affected if the data is placed on the target side, but they show greater resilience against such flaws in low-resource languages, which are less represented during pre-training.
@@ 2.2 Superficial alignment hypothesis
Zhou et al. (2023) claim that a model’s knowledge and capabilities are acquired almost entirely during pre-training, and the effect of alignment tuning might be “superficial”, in that it teaches the model the format for interacting with users. This idea is further supported by recent works (Lin et al., 2024; Ghosh et al., 2024). However, to what extent this applies to multilingual translation in LLMs is little known. To bridge this gap, we conduct a series of controlled experiments on fine-tuning LLMs for translation, complementing previous research across three dimensions. First, we study the parallel data efficiency in the era of LLMs, aiming to determine the minimum data needed for effective model alignment to the translation task. Next, we explore the scope of alignment by probing whether aligning one translation direction influences other directions. Finally, we investigate how synthesized fine-tuning data quality impacts the LLMs’ behaviour in generating translations.
@@ 3.1 Experimental setup
By default, we take the test sets from WMT17 to WMT20 as our parallel training data (Bojar et al., 2017, 2018; Barrault et al., 2019, 2020); we also use the development sets in WMT21 (Akhbardeh et al., 2021) for training if a language pair of interest is not available in earlier years. The specific training data configurations will be detailed in the subsequent sections. The test sets from WMT21 are used for validation. Detailed data statistics can be found in Section F.1. The LLM we use for SFT is the base version of Llama-2 7B (Touvron et al., 2023). When performing SFT, we use a learning rate of 5e-6, an effective batch size of 64, and a linear learning rate scheduling with a warmup ratio of 0.1. We select the model checkpoint based on COMET scores on the validation sets.222In our preliminary experiments, we found that validation perplexity has a relatively weak correlation with COMET scores measured on the validation set, similar to earlier findings (Ouyang et al., 2022). To form the model input for SFT, we feed the source sentence into the Alpaca prompt template (Taori et al., 2023), supplementing it with a translation instruction that is randomly selected from a pool of 31 diverse instructions. Refer to Table 4 in the appendix for a complete list of templates.
@@
We primarily evaluate the models on the WMT22 test sets (Kocmi et al., 2022) covering 11 translation directions: encs, ende, enjp, enru, enzh, and enhr.333Language codes: cs=Czech, de=German, hr=Croatian, jp=Japanese, ru=Russian, zh=Chinese. “” means that both translation directions are covered. Note that only enhr is available in WMT22 but not hren. Languages in these 11 directions are explicitly included in Llama-2’s pre-training corpus. In Section 3.4, we extend our evaluation to translation directions involving medium and low resource languages: Icelandic and Hausa (i.e., enis, enha), which comes from WMT21’s test set. At inference time, a fixed translation instruction is applied (Table 4 row 1). We use beam search with a beam size of 4 for generation, as our preliminary results indicate that it offers better translation quality than sampling-based generation, an observation consistent with recent works Jiao et al. (2023); Zeng et al. (2024). The maximum generation length is set to 256 tokens. We used a reference-based COMET22 checkpoint444Specifically, COMET is reported on a scale of 0 to 100 as opposed to its raw 0 to 1 range. (Rei et al., 2020) and BLEU (Papineni et al., 2002) as the evaluation metrics. See Section F.3 for detailed software configurations.
3.2 How much SFT data enables LLMs to translate?
@@
We fine-tune Llama-2 7B using different numbers of training samples and evaluate the multilingual translation performance of the resulting models. We collect training data covering 10 translation directions: en{cs, de, jp, ru, zh}. The training data sourced from WMT17-20 contains a total of 74,623 parallel examples. Note that the training samples across translation directions are not evenly distributed. To create training sets of varying sizes, we subsample the original data into subsets that are powers of 2, starting from 16 () and ending with 4096 (); larger subsets always contain smaller ones. To ensure balanced language representation in our subsets, we distribute samples as evenly as possible among the language pairs.555For example, the data size distribution for our 32-example training set is .
We refer to the fine-tuned model as SFT-MT. Considering LLMs can also perform translation through prompting, we compare SFT-MT with 1- and 3-shot in-context learning (ICL), denoted as ICL-MT. For ICL, we randomly select demonstrations from the training set in the test direction for each test sentence. We do not consider Llama-2’s zero-shot performance because, although it sometimes produces acceptable translations in the beginning, it often continues generating, which makes it difficult to accurately estimate its performance. Lastly, since LLMs fine-tuned on diverse tasks also serve as strong translation systems (Zhu et al., 2024), we compare our models with open-source general-purpose instruction-tuned LLMs, which we denote as IT-LLM. These include Vicuna-v1.5-7b (Chiang et al., 2023), Mistral-7b-Instruct (Jiang et al., 2023), and Llama-2-7b-chat (Touvron et al., 2023).666lmsys/vicuna-7b-v1.5, Mistral-7B-Instruct-v0.1, and meta-llama/Llama-2-7b-chat-hf.
@@ Results.
Figure 1 illustrates the effect of varying training sizes on translation performance. In both 1- and 3-shot cases, ICL-MT underperforms IT-LLM baselines like Llama-2-7b-chat despite sharing the same foundation model, indicating that a few in-context demonstrations may not effectively align Llama-2 for translation.
However, performance significantly improves when Llama-2 is fine-tuned with just 16 samples. With further increases in the training size to 32 samples, Llama-2 performs on par with or surpasses all three IT-LLM baselines in both COMET and BLEU metrics. This suggests that a handful of high-quality parallel data can effectively specialize the model into a performant translation system. Increasing parallel data further boosts performance, though with diminishing returns: the COMET score rises by an average of 2 points when expanding from 32 to 1024 samples, but only by 0.5 points when increasing further from 1024 to 75K samples (full training set). Given that it is unlikely that these 32 training samples “teach” Llama-2 new translation skills, this shows strong evidence that superficial alignment applies to MT. We observe a similar trend in Mistral-7B and Llama-2-13B. Refer to Appendix A for their performance across varying data sizes. In summary, effective translation alignment begins with minimal training data, revealing less is good alignment and more is better with diminishing gains.
@@ 3.3 Do we need to include all directions?
In the preceding section, we follow the traditional practice in multilingual MT by including multiple translation directions during training. However, the observation that only a few dozen examples make Llama-2 translate well leads us to reconsider the necessity of including samples from all directions of interest. Specifically, will training on just a single translation direction be sufficient to help LLMs perform multilingual translation?
@@
Figure 2 demonstrates the normalized performance of Llama-2 when fine-tuned in various single directions. Remarkably, training with just one direction enables Llama-2 to translate between multiple languages. For instance, after fine-tuning on deen or zhen, the model can translate from all considered languages to English, scoring at least of the original COMET scores for training on all directions. Similarly, the model fine-tuned on ende, enzh, frde or defr also demonstrates only a slight performance decline when translating from English.
Notable declines are observed in two scenarios: (1) trained to translate to English and evaluated on translating to non-English; and (2) trained to translate to non-English and evaluated on translating to English.777Analysis of model outputs reveals that they often merely echo the source sentence, ignoring the translation instruction. Of these two scenarios, scenario 1 exhibits a much larger performance drop. The fact that both scenarios involve a mismatch between using English and non-English suggests that Llama-2, as an English-centric LLM, may process English differently compared to other languages. When fine-tuned for English generation, the model may misinterpret the task as only generating in English. Generalization among non-English languages is much easier than generalization between English and non-English languages, as evidenced by the negligible performance drop when fine-tuning and testing on two vastly different language pairs such as defr and enzh. Overall, the findings suggest that SFT in one translation direction effectively enables the many directions, though avoiding misinterpretation is crucial.
@@
We also provide results of performing ICL with misaligned translation directions between demonstration and test in Table 1. It can be seen that misaligned demonstrations significantly degrade translation performance, with 3-shot be often worse than 1-shot. We observe that the model may output Chinese characters, emojis, time, etc., but no clear error patterns are observed. This contrasts sharply with findings from SFT: while SFT can recognize the format of translation, ICL requires language-aligned demonstrations.
@@
Figure 3 presents a joint evaluation of size and translation direction. For small training sizes, covering diverse translation directions in training proves to be beneficial. However, the benefits of such diversity level off as the training size increases. With a training size of 1024, models trained exclusively on two directions, ende, perform on par with those trained on all directions.
3.4 Can alignment be achieved for unseen languages?
Previous sections focus on translation directions involving languages explicitly included in Llama-2’s pre-training corpus. We now extend our investigation to languages that do not have an identified presence of over 0.005% in the pre-training data (c.f. Touvron et al., 2023, p22), referred to as unseen languages. Here we seek answers to two questions: (1) Can we effectively make Llama-2 translate both from and to unseen languages by fine-tuning it with a small amount of data? (2) How well can this fine-tuned model translate from and to languages seen in Llama?
@@ Results.
The results are presented in Figure 4. It can be seen that fine-tuning on Icelandic and Hausa enhances a model’s translation quality on these languages compared to the control setup, yet the gains are modest. We observe that Llama-2 manages to produce tokens in these languages, however, the translations often largely deviate from the original meanings. This suggests that it is difficult to teach models new translation directions via SFT with limited data. Interestingly, we find fine-tuning on Icelandic or Hausa does not hinder Llama-2’s ability to translate from and to all seen languages, maintaining performance levels comparable to the control scenario with ende. Based on these results, we propose a complement to the superficial alignment hypothesis in MT: LLMs may learn the essence of the translation task without requiring input-output mappings in languages it “understands” well.
| Source | Ref./Data config. | Model output |
@@
| sehr ärgerlich. | literal | The find I honest said very annoying. |
@@
| literal | in order to avoid again happen such thing. | |
@@
3.5 Can we use synthesized data?
We have observed that LLMs quickly recognize the translation task with minimal high-quality, manually curated data, but what if the quality of the training data is subpar? This situation may occur, for example when parallel data is web-crawled or machine-generated. Can LLMs still adapt to the translation task or will they overfit to the imperfections in lower-quality data, leading to degraded translation performance?
Setup.
We replace either the source or target sentences in the original training set with lower-quality synthesized ones. We try two types of data synthesis: one by translating entire sentences on the other side and another by concatenating word-to-word translations. Pleasingly, these correspond to back-translation (Sennrich et al., 2016) using translation engines or bilingual word dictionaries which are practical at different levels of resource availability. Specifically, we use the OPUS-MT suite (Tiedemann and Thottingal, 2020) to translate from English to a target non-English language.888E.g. for deen, the process is run in ende with the created data reversed, hence the translated content is on the source side. Checkpoints are available on Hugging Face: Helsinki-NLP/opus-mt-en-${trg}. For word-level translation, we translate each space-delimited source word by feeding it into the MT model one at a time. Naturally, the synthesized versions introduce translation errors, adding “noise” to the training process. We investigate the impact of such noise in four translation directions: ende′, de′en, enha′, and ha′en, where the prime (′) notation denotes the side that is created using translation (noised). We consider two training sizes: 32 and 1024. In this section, our evaluation focuses on the 11 translation directions described in Section 3.1. Note that although Hausa is included in the current training setup, translation directions involving Hausa are excluded from our evaluation—because performance is sub-par for unseen languages as demonstrated in Section 3.4.
Results.
According to Figure 5, it can be seen that both types of data synthesis generally cause a drop in performance. However, The degree of degradation significantly varies depending on whether the noise appears on the source or target side of the translation as well as the language. Specifically, when noise is introduced to the target side, models fine-tuned on ende′ and enha′ translations exhibit a sharp decline in performance. The impact of word noise is more severe than that of sentence noise. In the case of ende′, word-level synthesis causes the model to largely degenerate, leading to literal translations across many test cases across translation directions. An example of this behaviour is presented in Section 3.4. In contrast, the performance drop caused by word noise is less pronounced with enha′, particularly when evaluated on enX.
Summarizing the observations, Llama-2 is much more robust against the noise introduced in Hausa, likely because it has limited familiarity with the language, making it more difficult to detect and imitate imperfections present in the training data. As a result, Llama-2 tends to just recognize the essence of the translation task instead of overfitting to the biases present in low-quality data. In contrast, with German, Llama-2’s understanding leads to a misinterpretation of the training objectives, such as fitting the word-level noise with a directive for literal translations. Overall, LLMs may quickly fit translation imperfections in the training data, especially for seen languages; the resulting performance drop may be observable with just 32 training samples.